OpenStudy (anonymous):

solve the seperable equation: x(dv/dx)= (1-4v^2)/3v

OpenStudy (anonymous):

v=0, i√6/2,-i√6/2

OpenStudy (unklerhaukus):

v=0, x(dv/dx) = 1/0 !

OpenStudy (anonymous):

thank you! i just dont know how you reached that answer

OpenStudy (anonymous):

x(dv/dx)= (1-4v^2)/3v [3v/(1-4v^2)] dv =dx/x let u=1-4v^2, du=-8vdv -(3/8)du/u=dx/x -3/8 lnu =lnx +lnc1 -3/8 ln(1-4v^2) =lnx +lnc1 ln(1-4v^2)^(-3/8) =lnx +lnc1 (1-4v^2)^(-3/8) =xc1 1/[(1-4v^2)^(3/8)] =xc1 or 1/c1 = x(1-4v^2)^(3/8) C= x(1-4v^2)^(3/8)